# Proton Decay and Gauge Coupling Unification in an Extended SU(5) GUT with 45-Dimensional Higgs

### v2

@@ 2 SU(5) GUT with 45 dimensional Higgs
$+(Y_{6}^{DU})_{ij}\,\bar{d}_{Rai}\,(\eta^{A})^{ab}S_{6}^{A}\,u_{Rbj}^{C}+\frac{(Y_{6}^{QQ})_{ij}}{2}\,\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{\,C\,a\alpha}(\eta^{A})_{ab}\,S_{6}^{A*}\,q_{Lj}^{b\beta}$
$+(Y_{8}^{UQ})_{ij}\epsilon_{\alpha\beta}\,\bar{u}_{Rai}\,(\lambda^{A})^{a}{}_{b}\,S_{8}^{A\alpha}\,q_{Lj}^{b\beta}+(Y_{8}^{DQ})_{ij}\,\bar{d}_{Rai}\,(\lambda^{A})^{a}{}_{b}\,S_{8\alpha}^{A*}\,q_{Lj}^{b\alpha}+\mathrm{h.c.}\,,$ (6)
@@ 3 Gauge Coupling Unification
$\alpha_{1}^{-1}(M_{Z})=\alpha_{1}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $\frac{41}{10}\log\frac{M_{Z}}{\Lambda}+\frac{1}{15}\log\frac{M_{H_{C}}}{\Lambda}-\frac{35}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{10}\log\frac{M_{H^{(45)}}}{\Lambda}$
@@
$\alpha_{2}^{-1}(M_{Z})=\alpha_{2}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $-\frac{19}{6}\log\frac{M_{Z}}{\Lambda}-\frac{21}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{3}\log\frac{M_{\Sigma}}{\Lambda}+\frac{1}{6}\log\frac{M_{S_{1}}}{\Lambda}$
@@
$\alpha_{3}^{-1}(M_{Z})=\alpha_{3}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $-7\log\frac{M_{Z}}{\Lambda}-\frac{1}{6}\log\frac{M_{H_{C}}}{\Lambda}-7\log\frac{M_{XY}}{\Lambda}-\frac{1}{2}\log\frac{M_{\Sigma}}{\Lambda}$
@@
$-5\alpha_{1}^{-1}(M_{Z})+3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl{[}$ $-42\log\frac{M_{XY}}{M_{Z}}-2\log\frac{M_{\Sigma}}{M_{Z}}+6\log\frac{M_{S_{3}}}{\Lambda}$
@@
$M_{S_{8}}=3\times 10^{4}\mathrm{GeV},\quad M_{S_{3}}=5\times 10^{7}\mathrm{GeV},\quad M_{S_{1}}=1.8\times 10^{12}~{}\mathrm{GeV}.$ (16)
@@ 4 Proton decay
$\lambda$ $A$ $\overline{\rho}$ $\overline{\eta}$

### v3

@@ 2 SU(5) GUT with 45 dimensional Higgs
$+(Y_{6}^{DU})_{ij}\,\bar{d}_{Rai}\,(\eta_{A})^{ab}S_{6}^{A}\,u_{Rbj}^{C}+\frac{(Y_{6}^{QQ})_{ij}}{2}\,\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{\,C\,a\alpha}(\eta_{A})_{ab}\,S_{6}^{A*}\,q_{Lj}^{b\beta}$
$+(Y_{8}^{UQ})_{ij}\epsilon_{\alpha\beta}\,\bar{u}_{Rai}\,(\lambda_{A})^{a}{}_{b}\,S_{8}^{A\alpha}\,q_{Lj}^{b\beta}+(Y_{8}^{DQ})_{ij}\,\bar{d}_{Rai}\,(\lambda_{A})^{a}{}_{b}\,S_{8\alpha}^{A*}\,q_{Lj}^{b\alpha}+\mathrm{h.c.}\,$, (6)
@@ 3 Gauge Coupling Unification
$\alpha_{1}^{-1}(M_{Z})=\alpha_{1}^{-1}(\Lambda)+\frac{1}{2\pi}\biggl[$ $\frac{41}{10}\log\frac{M_{Z}}{\Lambda}+\frac{1}{15}\log\frac{M_{H_{C}}}{\Lambda}-\frac{35}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{10}\log\frac{M_{H^{(45)}}}{\Lambda}$
@@
$\alpha_{2}^{-1}(M_{Z})=\alpha_{2}^{-1}(\Lambda)+\frac{1}{2\pi}\biggl[$ $-\frac{19}{6}\log\frac{M_{Z}}{\Lambda}-\frac{21}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{3}\log\frac{M_{\Sigma}}{\Lambda}+\frac{1}{6}\log\frac{M_{S_{1}}}{\Lambda}$
@@
$\alpha_{3}^{-1}(M_{Z})=\alpha_{3}^{-1}(\Lambda)+\frac{1}{2\pi}\biggl[$ $-7\log\frac{M_{Z}}{\Lambda}-\frac{1}{6}\log\frac{M_{H_{C}}}{\Lambda}-7\log\frac{M_{XY}}{\Lambda}-\frac{1}{2}\log\frac{M_{\Sigma}}{\Lambda}$
@@
$-5\alpha_{1}^{-1}(M_{Z})+3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl[$ $42\log\frac{M_{XY}}{M_{Z}}+2\log\frac{M_{\Sigma}}{M_{Z}}+6\log\frac{M_{S_{3}}}{\Lambda}$
@@
$M_{S_{8}}=3\times 10^{4}~\mathrm{GeV},\quad M_{S_{3}}=5\times 10^{7}~\mathrm{GeV},\quad M_{S_{1}}=1.8\times 10^{12}~\mathrm{GeV}.$ (16)
@@ 4 Proton decay
$\lambda$ AA $\overline{\rho}$ $\overline{\eta}$